# Gradient-based optimization of exact stochastic kinetic models

### v1

@@ Abstract
Stochastic kinetic models describe systems across biology, chemistry, and physics where discrete events and small populations render deterministic approximations inadequate. Parameter inference and inverse design in these systems require optimizing over trajectories generated by the Stochastic Simulation Algorithm, but the discrete reaction events involved are inherently non-differentiable. We present an approach based on straight-through Gumbel-Softmax estimation that maintains exact stochastic simulations in the forward pass while approximating gradients through a continuous relaxation applied only in the backward pass. We demonstrate robust performance on parameter inference in stochastic gene expression, accurately recovering kinetic rates of telegraph promoter models from both moment statistics and full steady-state distributions across diverse and challenging parameter regimes. We further demonstrate the method’s applicability to inverse design problems in stochastic thermodynamics, characterizing Pareto-optimal trade-offs between non-equilibrium currents and entropy production. The ability to efficiently differentiate through exact stochastic simulations provides a foundation for systematic inference and rational design across the many domains governed by continuous-time Markov dynamics.
Introduction
@@
The Chemical Master Equation (CME) provides the mathematical framework for describing the time evolution of probability distributions over discrete system states in stochastic kinetic models [42]. Analytical solutions, however, are available only for restricted classes of reaction networks [36, 38]. For general systems, exact sample trajectories can be generated using Gillespie’s Stochastic Simulation Algorithm (SSA) [10], which produces realizations of the underlying continuous-time Markov process. Parameter inference and inverse design in these systems therefore relies mostly on simulation-based approaches.
@@
We demonstrate the framework on two problem classes spanning biophysics and non-equilibrium statistical mechanics. First, we address parameter inference in stochastic gene expression, recovering kinetic rates of telegraph promoter models from both moment statistics and full steady-state probability distributions. Using differentiable histogram construction, we show that distribution-level optimization accurately recovers ground-truth parameters across kinetic regimes characteristic of prokaryotic transcription. Second, we apply the method to optimization problems in stochastic thermodynamics, maximizing non-equilibrium currents under constraints on entropy production and kinetic resources. We characterize Pareto-optimal trade-offs between current magnitude and dissipation, recovering known theoretical bounds. These applications illustrate the generality of the approach: any scalar objective computable from stochastic trajectories becomes accessible to efficient gradient-based optimization.
A Reparameterized Pathwise Gradient Estimator
@@
Figure 2: Moment matching inferences from synthetic data. Inference of parameters from mean and variance of the true underlying generative model at steady state, for 25 different parameter sets. Gray markers indicate the random initializations. (a) Two parameters of the telegraph model ($k_{\text{on}}$, $k_{\text{tx}}$) are simultaneously estimated. (b) Mean and standard deviation of the RNA levels at steady state for the true and optimized models. (c) Comparison between the estimated and true rates.
@@
Parameter Inference in Stochastic Gene Expression
We validate our inference framework on the telegraph promoter model, a canonical description of stochastic gene expression widely adopted in systems biology. First introduced by Peccoud and Ycart [30], the telegraph model has successfully explained experimental observations of transcriptional bursting across prokaryotic [12, 39] and eukaryotic systems [33, 40], and remains the basis for modern inference frameworks applied to single-cell RNA sequencing data [14].
The model describes a gene switching stochastically between two discrete promoter states: ON, permitting transcription, and OFF, where transcription is inhibited. The promoter transitions between states with rates $k_{\text{on}}$ and $k_{\text{off}}$. In the ON state, mRNA is produced at rate $k_{\text{tx}}$, while mRNA degradation occurs at rate $k_{\text{mdeg}}$. The system is represented by three reactions:
OFF $\xrightleftharpoons[k_{\text{off}}]{k_{\text{on}}}\text{ON}$
ON $\xrightarrow{k_{\text{tx}}}\text{ON}+\text{mRNA}$
mRNA $\xrightarrow{k_{\text{mdeg}}}\emptyset$
@@
Figure 3: Distribution matching inferences from synthetic data. Inference of parameters from the RNA histogram of the true underlying generative model at steady state, for 25 different parameter sets. Gray markers indicate the random initializations. (a) Three parameters of the telegraph model ($k_{\text{on}}$, $k_{\text{tx}}$, $k_{\text{mdeg}}$) are simultaneously estimated. (b) Comparison between the estimated and true transcription rate. (c) Relative errors in the estimation of the degradation rate $k_{\text{mdeg}}$ and the on-rate $k_{\text{on}}$ illustrate sloppy parameters directions. (d) Two representative examples comparing the target data distribution (black) and the distribution generated by the optimized model (red), showing agreement over several orders of magnitude.
Moment Matching
We first consider inferring two parameters of the telegraph model ($k_{\text{on}}$ and $k_{\text{tx}}$) while fixing the others to their ground-truth values (Fig. 2a). We initialize the free parameters randomly within a log-uniform range spanning $0.2\times,5\times]$ their true values. With only two free parameters, matching the mean and variance of the stationary distribution uniquely determines the solution.
@@
The optimized models accurately reproduce target moments (Fig. 2b), and recovers the true values of $k_{\text{on}}$ and $k_{\text{tx}}$ across all 25 parameter sets (Fig. 2c). We note that a previous study using automatic differentiation (but without the current methodology) on this class of problems achieved Pearson correlations of 0.68-0.74 using less challenging parameter regimes [35]. Full results are reported in Fig. S1.
Despite its apparent simplicity, this problem presents non-trivial optimization challenges. Although it admits a unique and analytically derivable minimum, the loss landscape is severely ill-conditioned around the global optimum, owing to loss landscapes featuring flat ridges corresponding to directions of low sensitivity [16]. This poses significant challenges to first-order optimizers, even when using adaptive methods such as Adam [21]. Achieving full convergence required deviating from standard hyperparameter settings, as detailed in Methods. The moment matching problem can be reformulated into a better-conditioned problem by reparameterizing to fit burst frequency $k_{\text{on}}k_{\text{off}}/(k_{\text{on}}+k_{\text{off}})$ and mean burst size $k_{\text{tx}}/k_{\text{off}}$, effectively preconditioning the landscape. However, identifying such reparameterizations is generally non-trivial, and we demonstrate here that our method is sufficiently robust to solve the original ill-conditioned problem directly.
Distribution Matching
We next consider inference from the full steady-state RNA copy-number distribution. The complete distribution contains substantially more information than moments alone and in principle can uniquely determine additional kinetic parameters. With three free parameters ($k_{\text{on}}$, $k_{\text{tx}}$, $k_{\text{mdeg}}$) and the remaining rate fixed to its ground-truth value (Fig. 3a), the optimization problem admits a unique minimum. In practice, however, the loss landscape exhibits pronounced near-degeneracies: coupled changes in promoter switching and mRNA degradation can produce nearly indistinguishable steady-state distributions, making the global optimum difficult to locate numerically.
@@
where $F_{\text{true}}$ and $F_{\text{sim}}$ denote the empirical CDFs of the target and simulated RNA counts, respectively. Our experiments showed that this objective is more robust than the more common information-theoretic alternatives: cross-entropy and KL divergence underweight low-probability bins and therefore neglects the tails, while symmetric divergences such as Jensen-Shannon become numerically unstable.
@@
Optimizing this loss requires differentiating through the simulated RNA histogram. For this we employ a triangular kernel density estimator with unit bandwidth. For integer-valued data, this produces histograms identical to standard hard binning, while providing well-defined gradients (Methods). Accurately resolving the distribution, particularly in the tails, requires a large number of simulations, but backpropagating through all of them would exceed even very loose memory constraints. We therefore compute distributional statistics from a large pool of forward-only simulations combined with a smaller set of gradient-tracked trajectories (Methods). This decouples the sample size needed for accurate distribution estimates from the memory cost of gradient computation, with the benefit of reducing gradient variance. Given the stiffness of this inverse problem, we adopt a multi-start protocol, running three independent optimizations from random initializations spanning $0.2\times,5\times]$ the ground-truth values and reporting the solution with the lowest loss.
@@
Across the 25 synthetic parameter sets, this procedure recovers $k_{\text{tx}}$ accurately (Fig. 3b). The remaining parameters exhibit correlated deviations (Fig. 3c): errors in $k_{\text{on}}$ and $k_{\text{mdeg}}$ are anticorrelated, reflecting a near-degenerate direction in the loss landscape where changes in switching kinetics can be partially compensated by changes in degradation. This pattern is consistent with the “sloppy parameter” structure characteristic of kinetic models [16]. Importantly, despite the difficulty of pinpointing the unique optimum along this degenerate direction, the optimized models faithfully reproduce the full target distribution over several orders of magnitude in probability (Fig. 3d; full results in Figs. S5S6). This demonstrates that ST-GS gradients provide sufficient signal to solve challenging distribution-level inverse problems while maintaining exact SSA dynamics throughout.
Figure 4: Optimal current-dissipation trade-off in a three-state ring. Optimization of cycle current $J$ under constraints on entropy production rate $\sigma$ and total kinetic budget $K_{\text{tot}}$. (a) Schematic of the three-state ring with forward ($k_{ij}^{+}$) and backward ($k_{ij}^{-}$) transition rates. (b) Optimized cycle current (blue circles) compared to the theoretical Pareto front (dashed line). Dotted line: asymptotic near-equilibrium scaling. Inset: Pareto front on linear axes. (c) Relative errors in $J$ (versus theory), $\sigma$, and $K_{\text{tot}}$ (versus target). Shaded band: $\pm 1\%$. (d) Optimized forward rates $k_{12}^{+}$, $k_{23}^{+}$, $k_{31}^{+}$ (light green symbols), their mean $\langle k^{+}\rangle$ (green diamonds), and backward rate $k^{-}$ (squares), compared to theoretical predictions (dashed and dotted lines).
Constrained Optimization in Nonequilibrium Stochastic Thermodynamics
We next apply our framework to an inverse design problem in stochastic thermodynamics [37]: characterizing the Pareto front of optimal trade-offs between steady-state current and entropy production in a driven system with fixed kinetic resources. This problem combines core ingredients of non-equilibrium thermodynamics (cycle currents, thermodynamic affinities, dissipation constraints) in a minimal setting that admits analytical solution, providing a rigorous benchmark.
We study a three-state ring [17] where a particle transitions cyclically among states $S_{1}\to S_{2}\to S_{3}\to S_{1}$, with each edge supporting both forward and backward transitions (Fig. 4a):
When detailed balance is broken, a net cycle current $J$ circulates through the ring. The thermodynamic cost of maintaining this current is the entropy production rate $\sigma=J\cdot\mathcal{A}$, where $\mathcal{A}=\sum_{i}\log(k_{i}^{+}/k_{i}^{-})$ is the cycle affinity [36]. We additionally impose a kinetic budget $K_{\text{tot}}=\sum_{i}(k_{i}^{+}+k_{i}^{-})$, which functions as a bound on the system’s dynamical activity [24].

### v2

@@ Abstract
Stochastic kinetic models describe systems across biology, chemistry, and physics where discrete events and small populations render deterministic approximations inadequate. Parameter inference and inverse design in these systems require optimizing over trajectories generated by the Stochastic Simulation Algorithm, but the discrete reaction events involved are inherently non-differentiable. We present an approach based on straight-through Gumbel-Softmax estimation that maintains exact stochastic simulations in the forward pass while approximating gradients through a continuous relaxation applied only in the backward pass. We demonstrate robust performance on parameter inference in stochastic gene expression, first recovering kinetic rates of telegraph promoter models from both moment statistics and full steady-state distributions across diverse and challenging synthetic parameter regimes, then inferring the kinetic parameters of a four-state promoter model from experimental single-molecule RNA timecourse measurements. We further apply the method to inverse design in stochastic thermodynamics, optimizing non-equilibrium currents in an interacting particle system under kinetic resource constraints and recovering known analytical bounds. The ability to efficiently differentiate through exact stochastic simulations provides a foundation for systematic scalable inference and rational design across the many domains governed by continuous-time Markov dynamics.
I Introduction
@@
The Chemical Master Equation (CME) provides the mathematical framework for describing the time evolution of probability distributions over discrete system states in stochastic kinetic models [49]. Analytical solutions, however, are available only for restricted classes of reaction networks [42, 44]. For general systems, exact sample trajectories can be generated using Gillespie’s Stochastic Simulation Algorithm (SSA) [11], which produces realizations of the underlying continuous-time Markov process. Parameter inference and inverse design in these systems therefore rely mostly on simulation-based approaches.
@@
We demonstrate the framework on two problem classes spanning biophysics and non-equilibrium statistical mechanics. For parameter inference in stochastic gene expression, we first validate the approach on synthetic data, recovering kinetic rates of telegraph promoter models from both moment statistics and full steady-state probability distributions. We then fit a four-state promoter model to experimental smFISH timecourse data, inferring eight kinetic parameters simultaneously from time-resolved RNA count distributions. For inverse design in stochastic thermodynamics, we optimize non-equilibrium currents in an interacting particle system under kinetic resource constraints, recovering known theoretical bounds. These applications illustrate the generality of the approach: any scalar objective computable from stochastic trajectories becomes accessible to efficient gradient-based optimization.
II Straight-Through Gumbel-Softmax Pathwise Gradient Estimator
@@
Figure 2: Moment matching inferences from synthetic data. Inference of parameters from the mean and variance of the true underlying generative model at steady state, for 25 different parameter sets. Gray markers indicate the random initializations. (a) Two parameters of the telegraph model ($k_{\mathrm{on}}$, $k_{\mathrm{tx}}$) are simultaneously estimated. (b) Mean and standard deviation of the RNA levels at steady state for the true and optimized models. (c) Comparison between the estimated and true rates.
@@
III Synthetic benchmarks with a telegraph promoter model
We first validate our inference framework on the telegraph promoter model, a canonical description of stochastic gene expression widely adopted in systems biology. First introduced by Peccoud and Ycart [36], the telegraph model has successfully explained experimental observations of transcriptional bursting across prokaryotic [13, 45] and eukaryotic systems [39, 47], and remains the basis for modern inference frameworks applied to single-cell RNA sequencing data [16].
The model describes a gene switching stochastically between two discrete promoter states: ON, permitting transcription, and OFF, where transcription is inhibited. The promoter transitions between states with rates $k_{\mathrm{on}}$ and $k_{\mathrm{off}}$. In the ON state, mRNA is produced at rate $k_{\mathrm{tx}}$, while mRNA degradation occurs at rate $k_{\mathrm{mdeg}}$. The system is represented by three reactions:
OFF $\xrightleftharpoons[k_{\mathrm{off}}]{k_{\mathrm{on}}}\text{ON}$
ON $\xrightarrow{k_{\mathrm{tx}}}\text{ON}+\text{mRNA}$
mRNA $\xrightarrow{k_{\mathrm{mdeg}}}\emptyset$
@@ III Synthetic benchmarks with a telegraph promoter model
Figure 3: Distribution matching inferences from synthetic data. Inference of parameters from the RNA histogram of the true underlying generative model at steady state, for 25 different parameter sets. Gray markers indicate the random initializations. (a) Three parameters of the telegraph model ($k_{\mathrm{on}}$, $k_{\mathrm{tx}}$, $k_{\mathrm{mdeg}}$) are simultaneously estimated. (b) Comparison between the estimated and true transcription rate. (c) Relative errors in the estimation of the degradation rate $k_{\mathrm{mdeg}}$ and the on-rate $k_{\mathrm{on}}$ illustrate sloppy parameter directions. (d) Two representative examples comparing the target data distribution (black) and the distribution generated by the optimized model (red), showing agreement over several orders of magnitude.
III.1 Moment Matching Loss
We first consider inferring two parameters of the telegraph model ($k_{\mathrm{on}}$ and $k_{\mathrm{tx}}$) while fixing the others to their ground-truth values (Fig. 2a). We initialize the free parameters randomly within a log-uniform range spanning $0.2\times,5\times]$ their true values. With only two free parameters, matching the mean and variance of the stationary distribution uniquely determines the solution.
@@ III.1 Moment Matching Loss
The optimized models accurately reproduce target moments (Fig. 2b), and recover the true values of $k_{\mathrm{on}}$ and $k_{\mathrm{tx}}$ across all 25 parameter sets (Fig. 2c). Full results are reported in Fig. S1. We note that a previous study using automatic differentiation (but without the current methodology) on this class of problems achieved Pearson correlations of 0.68-0.74 using less challenging parameter regimes [41].
Despite its apparent simplicity, this problem presents non-trivial optimization challenges. Although it admits a unique and analytically derivable minimum, the loss landscape is severely ill-conditioned around the global optimum, due to flat ridges corresponding to directions of low sensitivity [18]. This poses significant challenges to first-order optimizers, even when using adaptive methods such as Adam [24]. Achieving full convergence required deviating from standard hyperparameter settings, as detailed in Methods. The moment matching problem can be reformulated into a better-conditioned problem by reparameterizing to fit burst frequency $k_{\mathrm{on}}k_{\mathrm{off}}/(k_{\mathrm{on}}+k_{\mathrm{off}})$ and mean burst size $k_{\mathrm{tx}}/k_{\mathrm{off}}$, effectively preconditioning the landscape. However, identifying such reparameterizations is generally non-trivial, and we demonstrate here that our method is sufficiently robust to solve the original ill-conditioned problem directly.
Figure 4: Parameter inference from experimental single-molecule FISH timecourse. Inference of kinetic parameters for a four-state promoter model from single-molecule FISH measurements of STL1 osmotic stress response in yeast. (a) Schematic of the four-state promoter model: the promoter transitions between states 1–4 with rates $k_{ij}$, transcription occurs from the final state $S_{4}$ at rate $k_{\mathrm{tx}}$, and mRNA gets depleted at rate $k_{\mathrm{mdeg}}$. (b) Kullback-Leibler divergence between observed and simulated RNA distributions during optimization. (c) Convergence of kinetic rates over training iterations. (d) Comparison between observed RNA count distributions (black triangles: replicate average; colored circles: individual replicates) and the fitted model (red squares) at six of the eight fitted time points (selected for clarity) following osmotic stress, on a logarithmic probability scale.
III.2 Distribution Matching Loss
We next consider inference from the full steady-state RNA copy-number distribution. The complete distribution contains substantially more information than moments alone and in principle can uniquely determine additional kinetic parameters. With three free parameters ($k_{\mathrm{on}}$, $k_{\mathrm{tx}}$, $k_{\mathrm{mdeg}}$) and the remaining rate fixed to its ground-truth value (Fig. 3a), the optimization problem admits a unique minimum. In practice, however, the loss landscape exhibits pronounced near-degeneracies: coupled changes in promoter switching and mRNA degradation can produce nearly indistinguishable steady-state distributions, making the global optimum difficult to locate numerically.
@@ III.2 Distribution Matching Loss
where $F_{\text{true}}$ and $F_{\text{sim}}$ denote the empirical CDFs of the target and simulated RNA counts, respectively.
@@
Optimizing this loss requires differentiating through the simulated RNA histogram. For this we employ a triangular kernel density estimator with bandwidth $h=1.01$. For integer-valued data, this closely approximates standard hard binning while providing well-defined gradients (Methods). Accurately resolving the distribution, particularly in the tails, requires a large number of simulations, but backpropagating through all of them would exceed even very loose memory constraints. We therefore compute distributional statistics from a large pool of forward-only simulations combined with a smaller set of gradient-tracked trajectories (Methods). This decouples the sample size needed for accurate distribution estimates from the memory cost of gradient computation, with the benefit of reducing gradient variance. Given the stiffness of this inverse problem, we adopt a multi-start protocol, running three independent optimizations from random initializations spanning $0.2\times,5\times]$ the ground-truth values and reporting the solution with the lowest loss.
@@
Across the 25 synthetic parameter sets, this procedure recovers $k_{\mathrm{tx}}$ accurately (Fig. 3b). The remaining parameters exhibit correlated deviations (Fig. 3c): errors in $k_{\mathrm{on}}$ and $k_{\mathrm{mdeg}}$ are anticorrelated, reflecting a near-degenerate direction in the loss landscape where changes in switching kinetics can be partially compensated by changes in degradation. This pattern is consistent with the “sloppy parameter” structure characteristic of kinetic models [18]. Importantly, despite the difficulty of pinpointing the unique optimum along this degenerate direction, the optimized models faithfully reproduce the full target distribution over several orders of magnitude in probability (Fig. 3d; full results in Figs. S5S6). This demonstrates that ST-GS gradients provide sufficient signal to solve challenging distributional inverse problems while maintaining exact SSA dynamics throughout.
IV Parameter Inference from experimental gene expression timecourse distributions
Inferring kinetic parameters of stochastic gene expression models from single-cell measurements is a central challenge in quantitative biology, with growing interest in fitting such models at scale [15]. Because fitting the full RNA distribution shape—rather than summary statistics—is essential for producing mechanistic predictive models [31], inference methods must be both computationally efficient and expressive enough to handle complex dynamics. Differentiable stochastic simulation addresses these requirements naturally, enabling direct gradient-based optimization over full RNA count distributions without moment closures, state-space truncations, or likelihood-free approximations.
We demonstrate this approach by fitting a four-state promoter model to single-molecule fluorescence in situ hybridization (smFISH) data from the osmotic stress response of S. cerevisiae, using time-resolved measurements of STL1 nuclear mRNA following exposure to 0.4M NaCl [26]. This extends the two-state telegraph model to capture the multi-step activation process underlying transcriptional induction (Fig. 4a): the promoter transitions through a linear chain of four states with forward and reverse rates $k_{ij}$, accommodating effects such as chromatin remodeling and transcription factor binding. Transcription occurs exclusively from the final state $S_{4}$ at rate $k_{\mathrm{tx}}$, and nuclear mRNA is depleted at rate $k_{\mathrm{mdeg}}$ through degradation and nuclear export. This four-state topology is a simplified instance of the richer architecture originally developed to explain these data [33, 31], and has been independently supported by Bayesian model selection [23].
We restrict the analysis to the rising phase of the transcriptional response (0–20 min after stress), during which the upstream Hog1 kinase signal is active and rates can be approximated as constant. The data comprise smFISH measurements of nuclear STL1 mRNA counts at eight time points ($t=0,2,4,6,8,10,15,20$ min), pooled across three biological replicates. All cells are initialized in state $S_{1}$ and with zero mRNA molecules, reflecting the basal transcriptionally silent condition prior to osmotic stress. To match experimental count distributions we minimize the Kullback–Leibler divergence, a commonly used loss for distribution fitting, between observed and simulated RNA histograms simultaneously across all eight time points. Simulated distributions are constructed from a combined pool of gradient-tracked and baseline trajectories, as described in Distribution Matching. All eight kinetic parameters—the six inter-state transition rates $k_{ij}$, the transcription rate $k_{\mathrm{tx}}$, and the mRNA degradation rate $k_{\mathrm{mdeg}}$—are optimized simultaneously in log-space using the Adam optimizer with default hyperparameters, reflecting the standard choices we expect most practitioners to adopt.
The optimization converges within approximately 100 iterations (Fig. 4b), requiring less than five minutes of wall-clock time on a single NVIDIA A100 GPU. Kinetic rate constants separate into physically meaningful values (Fig. 4c), and the fitted model captures the time evolution of the full mRNA count distributions, reproducing both the dominant zero-count peak at early times and the gradual emergence of higher copy numbers as the transcriptional response develops (Fig. 4d; Figs. S8S11).
The fitted values (Table S1) are physically compatible with an induced yeast stress gene: the nuclear mRNA decay rate $k_{\mathrm{mdeg}}=0.095~\mathrm{min^{-1}}$ implies a $\sim 7.3$ min half-life, and the active-state production rate $k_{\mathrm{tx}}=0.434~\mathrm{min^{-1}}$ corresponds to one transcript every $\sim 2.3$ min while transcription-competent. These values are of the same order of magnitude as previous estimates for this system, though direct comparison is complicated by differences in model structure [31]. The promoter progression is strongly biased toward activation, with forward/reverse ratios increasing along the chain ($k_{12}/k_{21}\approx 3.50$, $k_{23}/k_{32}\approx 6.33$, $k_{34}/k_{43}\approx 9.13$), consistent with increasingly committed steps over the induction window.
Figure 5: Fundamental diagram of the Asymmetric Simple Exclusion Process. Optimization of particle current in the ASEP on a periodic lattice. (a) Schematic of the ASEP with $L$ sites and NN particles, each particle hops forward with rate $k_{i}^{+}$ and backward with rate $k^{-}$. (b) Current $J$ versus particle density $\rho=N/L$ for $L=10$ (blue circles) and $L=30$ (red squares), compared to finite-size theory (dotted line) and mean-field theory (dashed line). (c) Relative error in optimized current versus density, shaded band indicates $\pm 3\%$. (d) Coefficient of variation (CV) of optimized forward rates during training, showing convergence toward uniform rate allocation.
V Inverse design in nonequilibrium stochastic thermodynamics
We next apply our framework to an inverse design problem in stochastic thermodynamics [43]: optimizing steady-state current in an asymmetric simple exclusion process (ASEP) [46, 8] on a periodic lattice under a kinetic resource constraint. The ASEP couples exclusion interactions to nonequilibrium driving on a minimal many-body model with an analytically characterized optimum, providing a rigorous benchmark for gradient-based optimization of interacting particle systems.
We study a reversible ASEP ring with $L$ sites and hard-core exclusion, occupied by $N$ particles at density $\rho=N/L$ (Fig. 5a). Each bond (i,i+1)(i,i{+}1) supports both forward and backward hopping events:
$P_{i}+H_{i+1}$ $\xrightleftharpoons[k^{-}]{k_{i}^{+}}H_{i}+P_{i+1},$
where $P_{i}$ and $H_{i}$ denote a particle and a hole at site $i$, respectively. The forward rates $\{k_{i}^{+}\}_{i=1}^{L}$ are the design variables, while the backward rate $k^{-}$ is fixed and homogeneous across bonds, establishing a reference scale that removes the degeneracy between absolute rates and their ratios. Breaking detailed balance induces a net stationary particle current $J$ around the ring, with corresponding entropy production rate $\sigma=J\,\mathcal{A}$, where $\mathcal{A}=\sum_{i=1}^{L}\log(k_{i}^{+}/k^{-})$ is the cycle affinity.
The inverse design objective is to maximize $J$ subject to a fixed average forward hopping budget $\bar{k}^{+}=1/L\sum_{i}k_{i}^{+}$. This prevents the trivial strategy of uniformly increasing all rates and isolates the question of how to optimally distribute limited kinetic resources across bonds.